# Size-structured populations with growth fluctuations: Feynman–Kac formula and decoupling

### v1

@@ 6 Symmetric division and the mass-weighted phenotype density of path expectations
This section provides a more general interpretation of the tilted expectation equation (35), which holds when the size and growth phenotypes do not decouple, but division is symmetric. The argument is based on previously developed lineage representations of the population growth rate given in references [26, 37, 20], which we review in the following subsection.
@@ 6.1 The population tree argument and population distribution
Consider a population starting from one cell at time $t=0$. Suppose one selects a lineage from a tree by choosing either daughter cell at random at division. Let $K(t)$ denote the number of divisions along a lineage occurring before time $t$. Due to the symmetry of divisions, the distribution of $K(t)$ should be identical for each lineage in a growing tree. Now fix the number of generations $k$ (assuming the first cell is of generation zero for simplicity). The population tree contains $2^{k}$ $k$-th generational cells, each of which can be associated with one of $2^{k}$ lineages that are unique at the $k$-th generation. These lineages are indexed by $j\in\{1,\dots 2^{k}\}$ and have a division count function $K_{j}(t)$. The number of cells in generation $k$ at time $t$ can be obtained by counting the number of lineages $j$ that have $K_{j}(t)=k$, and hence it is given by the sum $\sum_{j=1}^{2^{k}}1_{\{K_{j}(t)=k\}}$. Although the different lineage division counts are highly dependent on each other, the expected number of $k$-th generational cells $n(k,t)$ can be computed via
$n(k,t)=\mathbb{E}_{\rm l}\left[\sum_{j=1}^{2^{k}}1_{\{K_{j}(t)=k\}}\right]=\sum_{j=1}^{2^{k}}\mathbb{P}\left(K_{j}(t)=k\right)=2^{k}\mathbb{P}\left(K(t)=k\right)$ (46)
The expected total number of cells at time $t$ is therefore
$n(t)=\sum_{k=0}^{\infty}n(k,t)=\sum_{k=0}^{\infty}2^{k}\mathbb{P}(K(t)=k)=\mathbb{E}_{\rm l}^{\rm path}\left[2^{K(t)}\right]$ (47)
This automatically gives a lineage representation of the long-term growth rate [26].
We make a similar argument to obtain expectations with respect to the population phenotype distribution, $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]$. If we say a population at time $t$ consists of cells $i,\dots,N(t)$, with $N(t)$ being a random variable, then the population distribution can be given as
$\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]=\frac{1}{n(t)}\mathbb{E}\left[\sum_{i=1}^{N(t)}f(X_{i}(t))\right].$ (48)
For any population, the summation over $\{1,\dots,N(t)\}$ can be broken down into a summation over all possible cells in the entire population tree, each cell given by its generation $k\in\{1,\dots\}$ and lineage index $j\in\{1,\dots,2^{k}\}$, where cells are only counted if they are actually present in the population at time $t$. This condition is satisfied only if $K_{j}(t)=k$. Hence, we write
$\mathbb{E}_{\rm p}\left[f({\boldsymbol{X}}(t))\right]=\frac{1}{n(t)}\mathbb{E}\left[\sum_{k=0}^{\infty}\sum_{j=1}^{2^{k}}f({\boldsymbol{X}}_{j}(t))1_{\{K_{j}(t)=k\}}\right].$ (49)
Given the generation of a cell $k$, the distribution of $X_{j}(t)$ and $K_{j}(t)$ are the same for all cells $j$, hence we may drop the index $j$ and evaluate the sum as
$\frac{1}{n(t)}\sum_{k=0}^{\infty}2^{k}\mathbb{E}_{\rm l}^{\rm path}\left[f({\boldsymbol{X}}_{j}(t))1_{\{K(t)=k\}}\right]=\frac{1}{n(t)}\mathbb{E}_{\rm l}^{\rm path}\left[f({\boldsymbol{X}}(t))2^{K(t)}\right].$ (50)
By using $n(t)=\mathbb{E}_{l}[2^{K(t)}]$, we obtain
@@
$\mathbb{E}_{\rm p}\left[f({\boldsymbol{X}}(t))\right]=\frac{\mathbb{E}_{\rm l}^{\rm path}\left[f({\boldsymbol{X}}(t))2^{K(t)}\right]}{\mathbb{E}_{\rm l}^{\rm path}\left[2^{K(t)}\right]}.$ (51)
@@
This is a general formula that holds irrespective of the division and growth dynamics. For example, even if $\mathcal{L}$ depends on the size variable, we can still obtain population expectations in this way. Next, we will show more generally, in the case of SD of section 3, this same argument can be repeated with mass instead of division numbers to derive equation (41).
@@ 6.2 Connecting Feynman–Kac expectation to mass-weighted statistics
$\mathbb{E}_{\rm l}[f({\boldsymbol{X}}(t))e^{T(t)}]=\mathbb{E}_{\rm l}[f({\boldsymbol{X}}(t))e^{Y_{c}(t)}2^{K(t)}].$ (53)
In particular, $\mathbb{E}_{\rm l}[e^{Y_{c}(t)}2^{K(t)}]$ is the expected total biomass of the population at time $t$.
Just as we have a number density $n({\boldsymbol{x}},t)$, we can introduce a mass-weighted population density of the phenotype $\boldsymbol{X}$}
@@
With this notation, the identity above reads
$\mathbb{E}_{\rm l}[f({\boldsymbol{X}}(t))\,e^{Y_{c}(t)}2^{K(t)}]=\int f({\boldsymbol{x}})\,m({\boldsymbol{x}},t)\,d{\boldsymbol{x}}.$ (55)
Then, in general, tilted expectations of the form given in equation (35) can be understood as averages with respect to this size-weighted density, and the finite-time version of equation (41) can be interpreted as the population mass-weighted expected value of $f({\boldsymbol{X}})$. Asymptotically, only when we have the decoupling condition, we can pull out the expected value of $e^{Y_{c}(t)}$ from both the numerator and the denominator of equation (41), recovering the count-based population distribution given in equation (51). Nevertheless, we can still interpret equation (41) as the mass-weighted expected value, even in the absence of the decoupling, and even in finite time.
@@ 7.1 Setup
We will start with the lineage distribution and omit their subscript l. The boundary conditions in our model define a linear operator $B:\mathcal{M}\to\mathcal{M}^{b}$ which maps the joint density of growth phenotype, size, and initial size to the density at $y_{c}=y_{b}$. $B$ acts on a function $\rho\in\mathcal{M}$ according to
@@
This is simply the right-hand side of equation (8) written as a linear operator and divided by $\lambda({\boldsymbol{x}})$. We are using parentheses around $G\rho^{b}$ to make it clear that the input is a function of only $({\boldsymbol{x}},{\boldsymbol{y}})$, while the output is evaluated at $({\boldsymbol{x}},y_{b})$.
We next introduce an operation which takes functions in $\mathcal{M}^{b}$ and lifts them to $\mathcal{M}$. This operation is obtained from a $C_{0}$-semigroup $(\Pi(y))_{y\geq 0}$ on $\mathcal{M}$, meaning $\Pi(y)\rho_{b}$ is a continuous function from $\mathbb{R}_{\geq 0}$ to $\mathcal{M}^{b}$ for any $\rho_{b}\in\mathcal{M}^{b}$. This comes from solving the time-independent version of equation (6) (the PDE for $\rho_{\rm l}$) and treating size as the time variable. More concretely, for $\rho^{b}\in\mathcal{M}^{b}$ we define $G:\mathcal{M}^{b}\to\mathcal{M}$ by
@@
where $u({\boldsymbol{x}},y,y_{b})=(\Pi(y)\rho^{b})({\boldsymbol{x}},y_{b})$ is a solution to
@@
The equations for the steady-state lineage density and lineage birth density are normalized solutions to
$\rho_{\rm l}=GB\rho_{\rm l},\quad\rho_{\rm l}^{b}=BG\rho_{\rm l}^{b}.$ (60)
Notice that the second equation is nothing but a restatement of equation (8) in the long-time limit. We conjecture that, under reasonable assumptions on the model inputs, $GB$ and $BG$ have unique normalizable eigenfunctions corresponding to the maximum eigenvalue, which is $1$.
@@
Similar to equation (60), we then have
$\rho_{\rm p}=2\hat{G}_{\Lambda}B\rho_{\rm p},\quad\rho_{\rm p}^{b}=2B\hat{G}_{\Lambda}\rho_{\rm p}^{b}.$ (63)
These equations are a type of nonlinear eigenvalue problem which must be solved for both the density and $\Lambda$. For the dominant eigenvalue $\Lambda$, the operators $2\hat{G}_{\Lambda}B\rho_{\rm p}$ and $2B\hat{G}_{\Lambda}\rho_{\rm p}^{b}$ are generally not stochastic operators, but equations (63) still have unique normalizable solutions.
@@ 7.2 Model structure under conditions (Condition 1) and (Condition 3)
The subscripts remind us that these operators depend respectively on $\lambda$ and $\varphi$, hence incorporating the two factors in the division rate function separately.
Meanwhile, Condition (Condition 3) allows us to solve equation (59), because $\tilde{\mathcal{A}}$ becomes
@@
Since the second term, a multiplication operator, only depends on $\boldsymbol{y}$}. The operator exponential can be split into the exponential of $\mathcal{L}$ and a multiplication by a function $\psi$ depending on the size variables:
@@
In order to make the role of the eigenvalue problem equation (16) apparent, we introduce
@@
Observe that $\nu\in{\rm ker}(\mathcal{L}-(\Lambda+\lambda))$ if and only if $\nu$ solves equation (16), therefore $\hat{G}_{\Lambda}$ leaves solutions to equation (16) unchanged.
Just as we found for the lineage equations, the steady-state population size distribution for the size-only process is a fixed point of the operator $2\hat{\Psi}U$ where $\hat{\Psi}$ multiplies function in $\mathcal{M}_{y}^{b}$ by $\hat{\psi}(y_{c},y_{2})$.
@@ 7.3 Statement of proof of main result
We are now ready to prove the result in section 3. For completeness, we state them again here as a Theorem:
@@
Suppose Conditions (Condition 1) and (Condition 3) hold; that is, $(\mathcal{L},\lambda,\beta,h)=(\mathcal{L},\lambda,\lambda\varphi,ru)$. Additionally, assume the lineage and population distributions converge to unique time-invariant steady-states $\rho_{\rm l}$ and $\rho_{\rm p}$. Then, we have the following:
@@
  • In the case of WDl (resp. WDp), equation (15) holds for the lineage (resp. population) distribution.
@@ 7.4 Proof
The Theorem is a straightforward consequence of the formalism developed above. We only need to show the separable solutions do indeed solve equation (60) (resp. equation (63)) for the lineage (resp. population) distributions. We begin with the lineage distribution. If $\rho=\nu\cdot w$ and $\nu\in{\rm ker}(\mathcal{L})$ then $G(\nu\cdot w)=\nu\cdot\Psi w$. Hence if $w\in{\rm ker}(\Psi\circ U)$ and
$GB(\nu\cdot w)=G(R_{\lambda}\nu\cdot Uw)=\nu\cdot w$ (71)
when either $R_{\lambda}=I$ (SD) or $R_{\lambda}\nu=\nu$ (WDl). The proof for the population cases are identical, with ${\rm ker}(\mathcal{L})$ replaced by ${\rm ker}(\mathcal{L}-(\Lambda-\lambda))$.
@@
Regime Assumptions What factorizes How to compute
Strong decoupling (SD) $\beta=\lambda\,\varphi$, h=ru\,h=r\,u with $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})=\delta({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})$; $\boldsymbol{X}$} stationary ergodic Lineage and population: $\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\rm l}({\boldsymbol{x}})w_{\rm l}({\boldsymbol{y}})$ $\rho_{\rm p}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\rm p}({\boldsymbol{x}})w_{\rm p}({\boldsymbol{y}})$ $\nu_{\rm l}\in\ker\mathcal{L}$; $\nu_{\rm p}$ solves $(\mathcal{L}+\lambda)\nu_{\rm p}=\Lambda\nu_{\rm p}$; $w_{\rm l},w_{\rm p}$ are the size‑only solutions (section 4.1); $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]=\frac{\mathbb{E}_{\rm l}^{\rm path}\!\left[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}{\mathbb{E}_{\rm l}^{\rm path}\!\left[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}$.
Weak decoupling in lineage (WDl) $\beta=\lambda\,\varphi$, h=ru\,h=r\,u; $R_{\lambda}$ iteration converges and its fixed point lies in $\ker\mathcal{L}$ Lineage only: $\rho_{\rm l}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\rm l}({\boldsymbol{x}})w_{\rm l}({\boldsymbol{y}})$ $\nu_{\rm l}$ is the fixed point of $\nu\mapsto R_{\lambda}\nu$ (equation (18)); $w_{\rm l}$: size‑only; population generally does not factorize.
@@
Symmetric division, no decoupling $u$ symmetric (two equal daughters), no factorization assumed For any $t$: $\mathbb{E}_{\rm p}^{\rm mass}[f({\boldsymbol{X}}(t))]=\frac{\mathbb{E}_{\rm l}^{\rm path}\!\left[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}{\mathbb{E}_{\rm l}^{\rm path}\!\left[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}$; here $\mathbb{E}_{\rm p}^{\rm mass}$ is with respect to the mass‑weighted density $m({\boldsymbol{x}},t)$ (equation (54)).
General case $\mathbb{E}_{\rm p}\left[f({\boldsymbol{X}}(t))\right]=\frac{\mathbb{E}_{\rm l}\left[f({\boldsymbol{X}}(t))2^{K(t)}\right]}{\mathbb{E}_{\rm l}\left[2^{K(t)}\right]}$

### v2

@@ 6 Symmetric division and the mass-weighted phenotype density of path expectations
This section provides a more general interpretation of the tilted expectation equation (35), which holds when the size and growth phenotypes do not decouple, yet division is symmetric.The argument is based on previously developed lineage representations of population observables developed in [50] and later used to relate independent lineage statistics to population growth rates in  [39, 26, 24]. We review this in the following subsection.
@@ 6.1 The population tree argument and population distribution
Consider a population starting from one cell at time $t=0$ and select a lineage from a tree by randomly choosing either daughter cell at time $t$. In the following, we will consider the distributions of the $N(t)$ lineages in this tree.
For a given realization of the population tree, a lineage can be sampled in a forward manner by selecting a random cell at each division, which induces a distribution $\mathbb{P}_{f|{\rm tree}}$. The notation $f|{\rm tree}$ is to remind us that we are conditioning on the tree. If the lineages are labeled $1,\dots,N(t)$, the probability of sampling a particular lineage with label $i$ is $2^{-K_{i}}$, where $K_{i}$ is the number of divisions along lineage $i$.
A lineage can alternatively be sampled in a backward manner by randomly selecting a lineage from the $N(t)$ cells at time $t$. The probability of sampling a label $i$ under $\mathbb{P}_{b}$ is $1/N(t)$. Performing a change of measure relates the expectations for any observable $f({\boldsymbol{X}},{\boldsymbol{Y}})$ via
$\mathbb{E}_{b|{\rm tree}}[f({\boldsymbol{X}}_{i},{\boldsymbol{Y}}_{i})]=\frac{1}{N(t)}\mathbb{E}_{f|{\rm tree}}[f({\boldsymbol{X}}_{i},{\boldsymbol{Y}}_{i})2^{K_{i}}].$ (46)
Note that because we are conditioning on the tree structure, we can move $N(t)$ outside of the expectation.
Note that averaging the expectations of $\mathbb{P}_{f}$ and $\mathbb{P}_{b}$ over trees yields asymptotically $\mathbb{E}_{l}$ and $\mathbb{E}_{p}$ respectively. Hence
$\mathbb{E}_{p}[f({\boldsymbol{X}},{\boldsymbol{Y}})]$ $\sim\mathbb{E}[\mathbb{E}_{b|{\rm tree}}[f({\boldsymbol{X}}_{i},{\boldsymbol{Y}}_{i})]]=\mathbb{E}\left[\frac{1}{N(t)}\mathbb{E}_{f|{\rm tree}}[f({\boldsymbol{X}}_{i},{\boldsymbol{Y}}_{i})2^{K_{i}}]\right]$ (47)
where the outer expectation on the right is over the tree distribution. Note that if $f({\boldsymbol{X}},{\boldsymbol{Y}})=1$, then
$\mathbb{E}[\mathbb{E}_{f|{\rm tree}}[2^{K_{i}}]/N(t)]=1.$ (48)
Since $N(t)\sim e^{\Lambda t}$, this yields $\mathbb{E}_{f|{\rm tree}}[2^{K}_{i}]\sim Ce^{\Lambda t}$ where $C$ is a time-independent constant. The lineage expectation is related to the expectation on the right hand side of equation (47) via
@@
$\mathbb{E}\left[\frac{1}{N(t)}\mathbb{E}_{f|{\rm tree}}[f({\boldsymbol{X}}_{i},{\boldsymbol{Y}}_{i})2^{K_{i}}]\right]\sim e^{\Lambda t}\mathbb{E}_{l}^{\rm path}[f({\boldsymbol{X}}(t),{\boldsymbol{Y}}(t))2^{K(t)}]$ (49)
Combining these asymptotics we can relate expectations of the population observable to path expectations:
$\mathbb{E}_{\rm p}\left[f({\boldsymbol{X}}(t),{\boldsymbol{Y}}(t))\right]\sim\frac{\mathbb{E}_{\ell}^{\rm path}\left[f({\boldsymbol{X}}(t),{\boldsymbol{Y}}(t))2^{K(t)}\right]}{\mathbb{E}_{\ell}^{\rm path}\left[2^{K(t)}\right]}.$ (50)
@@
This is a general formula that holds irrespective of the division and growth dynamics. For example, even if $\mathcal{L}$ depends on the size variable, we can still obtain population expectations in this way.
@@ 6.2 Connecting Feynman–Kac expectation to mass-weighted statistics
$\mathbb{E}_{\ell}[f({\boldsymbol{X}}(t))e^{T(t)}]=\mathbb{E}_{\ell}[f({\boldsymbol{X}}(t))e^{Y_{c}(t)}2^{K(t)}].$ (52)
In particular, $\mathbb{E}_{\ell}[e^{Y_{c}(t)}2^{K(t)}]$ is the expected total biomass of the population at time $t$.
Using the number density, $n({\boldsymbol{x}},{\boldsymbol{y}},t)\sim e^{\Lambda t}\rho_{\rm p}({\boldsymbol{x}},{\boldsymbol{y}})$, we can introduce a mass-weighted population density of the phenotype $\boldsymbol{X}$} defined by
@@
The left hand side of equation (50) is, by definition, an expectation with respect to $\rho_{\rm p}\sim e^{-\Lambda t}n({\boldsymbol{x}},{\boldsymbol{y}},t)$. Meanwhile, equation (52) is of the form of the numerator on the right hand side of equation (50) when we make the replacement $f({\boldsymbol{X}},{\boldsymbol{Y}})\to f({\boldsymbol{X}})e^{Y_{c}}$. It follows that
$\mathbb{E}_{\ell}[f({\boldsymbol{X}}(t))\,e^{Y_{c}(t)}2^{K(t)}]\sim\int f({\boldsymbol{x}})\,m({\boldsymbol{x}},t)\,d{\boldsymbol{x}},$ (54)
Thus, in general, tilted expectations of the form given in equation (35) can be understood as averages with respect to this size-weighted density, and the finite-time version of equation (41) can be interpreted as the population mass-weighted expected value of $f({\boldsymbol{X}})$.
Only when we have the decoupling condition, we can pull out the expected value of $e^{Y_{c}(t)}$ from both the numerator and the denominator of equation (41), recovering the count-based population distribution given in equation (50). Nevertheless, we can still interpret equation (41) as the mass-weighted expected value, even in the absence of the decoupling, and even in finite time.
@@ 7.1 Setup
We will start with the lineage distribution and omit their subscript . The boundary conditions in our model define a linear operator $B:\mathcal{M}\to\mathcal{M}^{b}$ which maps the joint density of growth phenotype, size, and initial size to the density at $y_{c}=y_{b}$. $B$ acts on a function $\rho\in\mathcal{M}$ according to
@@
This is simply the right-hand side of equation (10) written as a linear operator and divided by $\lambda({\boldsymbol{x}})$. We are using parentheses around $B\rho^{b}$ to make it clear that the input is a function of $\boldsymbol{x}$} and the full size phenotype $\boldsymbol{y}$}, while the output is evaluated at $({\boldsymbol{x}},y_{b})$.
We next introduce an operation which takes functions in $\mathcal{M}^{b}$ and lifts them to $\mathcal{M}$. This operation is obtained from a $C_{0}$-semigroup $(\Pi(y))_{y\geq 0}$ on $\mathcal{M}$, meaning $\Pi(y)\rho_{b}$ is a continuous function from $\mathbb{R}_{\geq 0}$ to $\mathcal{M}^{b}$ for any $\rho_{b}\in\mathcal{M}^{b}$. This comes from solving the time-independent version of equation (7) (the PDE for $\rho_{\ell}$) and treating size as the time variable. More concretely, for $\rho^{b}\in\mathcal{M}^{b}$ we define $G:\mathcal{M}^{b}\to\mathcal{M}$ by
@@
where the function obtained by the application of $\Pi(y)$ to a function $\rho^{b}\in\mathcal{M}^{b}$, $u({\boldsymbol{x}},y,y_{b})=(\Pi(y)\rho^{b})({\boldsymbol{x}},y_{b})$, is a solution to
@@
The unique steady-state linear density satisfies the equation
$\rho_{\ell}=GB\rho_{\ell}.$ (59)
Note that because equation (59) is equivalent to the time-independent Kolmogorov equation for the lineage dynamics, it has a unique, normalizable solution.
@@
Similar to equation (59), we have
$\rho_{\rm p}=2\hat{G}_{\Lambda}B\rho_{\rm p}$ (62)
These equations are a type of nonlinear eigenvalue problem that must be solved for both the density and $\Lambda$.
@@ 7.2 Model structure under conditions (Condition 1) and (Condition 3)
The subscripts remind us that these operators depend, respectively, on $\lambda$ and $\varphi$, hence incorporating the two factors in the division rate function separately.
Meanwhile, Condition (Condition 1) allows us to factor the matrix exponential solution of  (58). Specifically, $\tilde{\mathcal{A}}$ becomes
@@
and the second term, a multiplication operator, only depends on $\boldsymbol{y}$}. As a result, the two operators are commutative and the operator exponential can be split into the exponential of $\mathcal{L}$ and a multiplication by a function $\psi$ depending on the size variables:
@@
In order to connect this to the eigenvalue problem, equation (17), we introduce
@@
Observe that $\nu\in{\rm ker}(\mathcal{L}-(\Lambda-\lambda))$ if and only if $\nu$ solves equation (17); therefore, $\hat{G}_{\Lambda}$ leaves solutions to equation (17) unchanged.
Just as we found for the lineage equations, the steady-state population size distribution for the size-only process is a fixed point of the operator $2\hat{\Psi}U$ where $\hat{\Psi}$ multiplies function in $\mathcal{M}_{y}^{b}$ by $\hat{\psi}(y_{c},y_{b})$.
@@ 7.3 Statement of proof of main result
We are now ready to prove the result in section 3. Our result includes an extension of the previously stated conditions, which will not be of biological interest but will help provide a more complete picture of the mathematical structure.
[…]
@@
Suppose Conditions (Condition 1) and (Condition 3) hold; that is, $(\mathcal{L},\lambda,\beta,h)=(\mathcal{L},\lambda,\lambda\varphi,ru)$. Additionally, assume the lineage and population distributions converge to unique time-invariant steady-states $\rho_{\ell}$ and $\rho_{\rm p}$. Then, we have the following:
@@
  • In the case of WD (resp. WDp), equation (16) holds for the lineage (resp. population) distribution.
@@ 7.4 Proof
The Theorem is a straightforward consequence of the formalism developed above. In particular, if the separable solutions do indeed solve equation (59) (resp. equation (62)) for the lineage (resp. population) distributions, then by assumption these are the unique, stable steady-state phenotype distributions.
Lineage distribution
The separable solution, if it exits, is $\rho=\nu\cdot w$, where $\nu\in{\rm ker}(\mathcal{L})$ and $w$ is the unique stationary densities of the size only process, i.e. $\Psi\circ Uw=w$. In the case of SD, $R_{\lambda}=I$ and therefore applying $G$ to $\nu\cdot w$ gives
$GB(\nu\cdot w)=G(\nu\cdot Uw)=\nu\cdot\Psi Uw=\nu\cdot w.$ (70)
In the case of WD, $R_{\lambda}\neq I$, but $R_{\lambda}\nu=\nu$ by assumption and therefore
$GB(\nu\cdot w)=G(\nu\cdot Uw)=R_{\lambda}\nu\cdot\Psi Uw=\nu\cdot w.$ (71)
Thus, in both SD and WD, $\rho=\nu\cdot w$ is a fixed point.
Population distribution
The proof for the population cases are identical, with the exception that instead of $\nu\in{\rm ker}(\mathcal{L})$ we have $\nu\in{\rm ker}(\mathcal{L}-(\Lambda-\lambda))$.
@@
Regime Condition What factorizes How to compute
Strong decoupling (SD) $\beta=\lambda\,\varphi$, h=ru\,h=r\,u with $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})=\delta({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})$; $\boldsymbol{X}$} stationary ergodic Lineage and population:  $\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\ell}({\boldsymbol{x}})w_{\ell}({\boldsymbol{y}})$ $\rho_{\rm p}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\rm p}({\boldsymbol{x}})w_{\rm p}({\boldsymbol{y}})$ $\nu_{\ell}\in\ker\mathcal{L}$; $\nu_{\rm p}$ solves $(\mathcal{L}+\lambda)\nu_{\rm p}=\Lambda\nu_{\rm p}$; $w_{\ell},w_{\rm p}$ are the size‑only solutions (section 4.1); $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]=\frac{\mathbb{E}_{\ell}^{\rm path}\!\left[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}{\mathbb{E}_{\ell}^{\rm path}\!\left[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}$.
Weak decoupling in lineage (WD) $\beta=\lambda\,\varphi$, h=ru\,h=r\,u; $R_{\lambda}$ iteration converges and its fixed point lies in $\ker\mathcal{L}$ Lineage only: $\rho_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})=\nu_{\ell}({\boldsymbol{x}})w_{\ell}({\boldsymbol{y}})$ $\nu_{\ell}$ is the fixed point of $\nu\mapsto R_{\lambda}\nu$ (equation (19)); $w_{\ell}$: size‑only; population generally does not factorize.
@@
Symmetric division, no decoupling $u$ symmetric (two equal daughters), no factorization assumed For any $t$: $\mathbb{E}_{\rm p}^{\rm mass}[f({\boldsymbol{X}}(t))]=\frac{\mathbb{E}_{\ell}^{\rm path}\!\left[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}{\mathbb{E}_{\ell}^{\rm path}\!\left[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}\right]}$; here $\mathbb{E}_{\rm p}^{\rm mass}$ is with respect to the mass‑weighted density $m({\boldsymbol{x}},t)$ (equation (53)).
General case $\mathbb{E}_{\rm p}\left[f({\boldsymbol{X}}(t))\right]=\frac{\mathbb{E}_{\ell}\left[f({\boldsymbol{X}}(t))2^{K(t)}\right]}{\mathbb{E}_{\ell}\left[2^{K(t)}\right]}$